OpenStudy (anonymous):

|2x+2|=14 solve for x

OpenStudy (anonymous):

there are two answers

OpenStudy (anonymous):

2x+2=positive or negative 14 so 1-2x+2=14 2x=12 x=6 _________- 2-2x+2=-14 2x=-16 x=-8

OpenStudy (anonymous):

the two answers are , (6,-8)